Please help me
Jack usually mows his lawn in 3 hours. Marilyn can mow the same yard in 7 hours. How much time would it take for them to mow the lawn together??

Answer by ewatrrr(24785)


Hi,
Jack usually mows his lawn in 3 hours. Marilyn can mow the same yard in 7 hours.
Let x represent the time it takes them together
1/3 + 1/7 = 1/x
Mulitply thru with 21x to clear the denominators
7x + 3x = 21
10x = 21
x = 2.1 hrs together